# Reference.bib

### v1

@@ 2.3.3. Examples for projective completions
Though the construction of the CMCX $\mathcal{T}$ and the completion $\hat{B}$ of $B$ given by Theorem 2.8 are in general unapproachable, there are still some accessible cases including the case when $B$ is a surface. In Section 7 we will review several such examples. We will show the main Theorem 2.8 gives completed period mappings in the category of projective varieties. These are cases in which the Conjecture 2.11 is true.
@@ 2.3.4. Algebraicity of weakly special subvarieties
Fix a weak Mumford-Tate domain $M$ with $\mathbb{H}:=\mathrm{Aut}(M)^{+}$. Let $\Gamma_{H}:=\Gamma\cap H_{\mathbb{Q}}$. A weakly special subvariety associated with the weak Hodge datum $(\mathbb{H},M)$ is an irreducible component of $\Phi^{-1}(\Gamma_{H}\backslash M)$ in $B$. It is an analytic subvariety of $B$.
According to the theory of functional transcendence and o-minimal geometry, since the period map is definable [BKT20], $\Phi^{-1}(\Gamma_{H}\backslash M)$ is a definable analytic subvariety of $B$ and hence must be algebraic by definable GAGA theorem [BBT22]. In particular, weakly special subvarieties are algebraic.
As an application of our main theorem, we give an alternative proof of the algebraicity result on weakly special subvarieties without using o-minimal methods. Our proof is based on the Hodge-theoretic interpretation of the completed period map $\overline{\Phi}$ provided by Theorem 2.8. We will show that weakly special subvarieties of $B$ can be extended to analytic subvarieties of $\overline{B}$ over which an extended period map $\overline{\Phi}$ exists. The boundary Hodge theoretic data are described by the period torsors defined by Kato-Usui in [KU08, Chap. 3].
3. Hodge theory backgrounds
@@ 3.1. Basic settings
(3.1) $\Phi:B\rightarrow\Gamma\backslash D$
where $D$ is the classifying space of polarized Hodge structures defined on the integral polarized lattice $(V_{\mathbb{Z}},Q)$ and $\check{D}$ be its compact dual. Moreover, denote $\mathbb{G}=\mathrm{Aut}(H,Q)$ and $\mathfrak{g}:=\mathrm{Lie}(G)$, then $\Gamma\leq G_{\mathbb{Z}}$ is the monodromy group of the local system $\mathbb{V}\rightarrow B$:
@@
In addition, we call $\Gamma\backslash D$ a Hodge variety associated to the Hodge datum $(\mathbb{G},D)$.
@@ 3.2. Nilpotent cones and nilpotent orbits
Suppose an open convex polyhedral cone $\sigma\subset\mathfrak{g}_{\mathbb{Q}}$ is generated by commuting nilpotent elements $\{N_{1},...,N_{n}\}$. In other words,
(3.2) $\sigma:=\{\sum_{j=1}^{n}x_{j}N_{j}\ |\ x_{1},...,x_{n}\in\mathbb{Q}_{>0}\}.$
@@
    $\mathrm{exp}(\sum_{j=1}^{n}z_{j}N_{j})F\in D$ for $z_{j}\in\mathbb{C}$ and $\mathfrak{Im}(z_{j})>>0$ for any $j$.
In this case we also call $\mathcal{Z}:=\exp(\sigma_{\mathbb{C}})F\subset\check{D}$ a nilpotent orbit , denoted as $\mathcal{Z}:=(\sigma,F)$. Denote $\bar{A}(\sigma)\subset\check{D}$ as the set of points $F\in\check{D}$ such that $(\sigma,F)$ is a nilpotent orbit, and $A(\sigma):=\exp(\sigma_{\mathbb{C}})\backslash\bar{A}(\sigma)$.
It is a classical result in Hodge theory that the notions of nilpotent orbits and limiting mixed Hodge structures are equivalent:
@@
[CK82] $(\sigma,F)$ gives a nilpotent orbit if and only if $(W(\sigma),F)$ is a mixed Hodge structure polarized by any $N\in\sigma$, where $W(\sigma)$ is the Jacobson-Morozov weight filtration associated to any $N\in\sigma$ [^2] We assume the weight filtration is centralized at the weight same as the pure Hodge structures in $D$..
@@
Any element $g\in\mathbb{G}$ acts on the set of nilpotent orbits by:
$g(\sigma,F)\rightarrow(\mathrm{Ad}_{g}\sigma,gF).$
@@ 3.2.1. Schmid’s nilpotent orbit theorem
For any $b\in\hat{B}$, choose a local neighborhood around $b\in U\subset\hat{B}$ satisfies $U\cap B\cong(\Delta^{*})^{k}\times\Delta^{n-k}$ on which we may consider the local monodromy operators $\{T_{i}\}_{1\leq i\leq k}$ around each boundary divisor and their logarithms $\{N_{i}:=\log(T_{i})\}_{1\leq i\leq k}$. One of the main ingredients is the following intepretation of Schmid’s nilpotent orbit theorem [Sch73]:
@@
The cone $\mathrm{Conv}(\{N_{i}\}_{1\leq i\leq k})\subset\mathfrak{g}_{\mathbb{Q}}$ is a nilpotent cone, and there is a nilpotent orbit $(\sigma_{b},F_{b})$ canonically associated to $b$ for some $F_{b}\in\check{D}$.
@@ 3.3. Kato–Nakayama–Usui’s construction
Let $\Sigma$ be a collection of rational nilpotent cones in $\mathfrak{g}_{\mathbb{Q}}$, and $|\Sigma|$ be its support in $\mathfrak{g}_{\mathbb{Q}}$. For $\sigma\in\Sigma$, set $\Gamma(\sigma):=\Gamma\cap\mathrm{exp}({\mathbb{Q}_{\geq 0}}\sigma)$. This is a semigroup. Denote $\Gamma_{\sigma}:=\Gamma_{\sigma}$ as its associated group.
@@
    $\overline{\sigma_{\mathbb{Q}}}=\mathbb{Q}_{\geq 0}<\mathrm{log}(\Gamma(\sigma))>$ for any $\sigma\in\Sigma$.
@@
    For any $\sigma,\tau\in\Sigma$, either $\sigma\cap\tau=\emptyset$ or $\sigma=\tau$;
@@
    For any $\sigma,\tau\in\Sigma$, suppose $\sigma\cap\tau$ is not empty, and there exists $F\in\check{D}$ such that both $(\sigma,F)$ and $(\tau,F)$ are nilpotent orbits, then $\sigma=\tau$.
@@
(3.7) $D_{\Sigma}:$ =$=\{(\sigma,Z)\ |\ \sigma\in\Sigma,Z\ \text{is a}\ \sigma-\text{nilpotent orbit}\ \}$
@@
The first main result of Kato–Nakayama–Usui’s theory is about the (partial) completion of the Hodge variety $\Gamma\backslash D$:
@@
    For any $\sigma\in\Sigma$, by regarding $\sigma$ as a fan containing all faces of $\sigma$, the canonical map $\Gamma_{\sigma}\backslash D_{\sigma}\rightarrow\Gamma\backslash D_{\Sigma}$ is a local homeomorphism.
We refer readers to [KU08, Chap. 3] for the literature of logarithmic manifolds. Some important properties will be reviewed in Section 8. When $D$ is Hermitian, one may choose the (weak) fan $\Sigma$ to be the rational polyhedral decompositions constructed by classical theories (e.g., [AMRT10], [Nam76]), then the Kato–Nakayama–Usui’s construction coincide with the toroidal compactifications $\overline{\Gamma\backslash D}^{\Sigma}$. In general, the space $\Gamma\backslash D_{\Sigma}$ does not admit a structure of complex varieties.
@@
[KU08, Chap. 2] Given $\Gamma\leq G_{\mathbb{Z}}$ and $\Sigma$ which is a (weak) fan strongly compatible with $\Gamma$. We say a $\mathbb{Z}$-PVHS $\mathbb{V}\rightarrow B$ and a partial completion $B^{\prime}$ of $B$ is compatible with $(\Gamma,\Sigma)$ if:
@@
If the $\mathbb{Z}$-PVHS $\mathbb{V}\rightarrow B$ can be extended to a $\mathbb{Z}$-LPVHS on $B^{\prime}$ compatible with $(\Gamma,\Sigma)$, then there exists a unique morphism
@@
extending $\Phi:B\rightarrow\Gamma\backslash D$ to $B^{\prime}$.
@@
If $\overline{\Phi}$ is proper, then $B^{\prime}\xrightarrow{\overline{\Phi}}\overline{\Phi}(B^{\prime})$ is a morphism in the category of complex algebraic spaces.
@@
This theorem is in [Usu06, Sec. 5]. Though the original theorem takes the assumption that $B^{\prime}$ is compact (for example, take $B^{\prime}=\hat{B}$), the proof only uses the properness of $\overline{\Phi}$ and hence applies to our case verbatim. ∎
@@ 4. The combinatorial monodromy complex (CMCX)
4.1. The combinatorial monodromy cone (CMC) and adjacent map
For any $I\in\mathcal{I}$, we may choose any point $b\in Z_{I}^{*}$, a neighborhood $b\in E_{b}\subset\hat{B}$ such that $E_{b}^{*}:=E_{b}\cap B$ is isomorphic to $(\Delta^{*})^{|I|}\times\Delta^{\mathrm{dim}(B)-|I|}$, and a nearby point $b^{\prime}\in E_{b}^{*}$. For every $i\in I$ there is a monodromy logarithm $N_{i}\in\mathrm{End}(V_{b^{\prime},\mathbb{Q}})\simeq\mathfrak{g}_{\mathbb{Q}}$ around the divisor $Z_{i}$. Let $\sigma_{I,b^{\prime}}:=\mathrm{Conv}\{N_{i},i\in I\}^{\circ}$ (or $\overline{\sigma_{I,b^{\prime}}}=\mathrm{Conv}\{N_{i},i\in I\}$). There exists a linear map:
@@ 4.1. The combinatorial monodromy cone (CMC) and adjacent map
It is clear that the closure of the monodromy cone $\overline{\sigma_{I,b^{\prime}}}$ is isomorphic to the image of $\mathbb{Q}_{\geq 0}^{|I|}$ under the quotient map $\mathbb{Q}^{|I|}\rightarrow\mathbb{Q}^{|I|}/\mathrm{Ker}(\lambda_{I})$. Its combinatorial structure is determined by the index set $I\subset\mathcal{I}$ and $\mathrm{Ker}(\lambda_{I})$. It does not depend on the choice of $b\in Z_{I}^{*}$, the neighborhood $E_{b}$ around $b$, or the nearby point $b^{\prime}$.
@@
The combinatorial monodromy cone (CMC) $\boldsymbol{\sigma}_{I}$ is defined to be the image of $\mathbb{Q}_{>0}^{|I|}$ under the map $\mathbb{Q}^{|I|}\rightarrow\mathbb{Q}^{|I|}/\mathrm{Ker}(\lambda_{I})$.
@@
Starting from this point, for every $I\in\mathcal{I}$ we will denote $\sigma_{I}$ or $\tau_{I}$ as the original monodromy cones as a subset of some $\mathrm{End}(V_{b,\mathbb{Q}})$, while the corresponding boldface characters $\boldsymbol{\sigma}_{I}$ or $\boldsymbol{\tau}_{I}$ as the corresponding CMC’s.
For any sub-index $I^{\prime}\subset I$, the CMC $\boldsymbol{\sigma}_{I^{\prime}}$ is determined by the space $\mathrm{Ker}(\lambda_{I^{\prime}})$, and $\boldsymbol{\sigma}_{I^{\prime}}\simeq\mathrm{Conv}\{N_{i},i\in{I^{\prime}}\}^{\circ}\subset\mathrm{End}(V_{b^{\prime},\mathbb{Q}})$. The set-theoretic inclusion $\{\epsilon_{i},i\in{I^{\prime}}\}\hookrightarrow\{\epsilon_{i},i\in I\}$ induces $\mathbb{Q}^{|{I^{\prime}}|}\hookrightarrow\mathbb{Q}^{|I|}$, and $\mathrm{Ker}(\lambda_{I^{\prime}})=\mathrm{Ker}(\lambda_{I})\cap\mathbb{Q}^{|{I^{\prime}}|}$. We have the commutative diagram:
@@
The map $\mathbb{Q}^{|{I^{\prime}}|}/\mathrm{Ker}(\lambda_{J})\hookrightarrow\mathbb{Q}^{|I|}/\mathrm{Ker}(\lambda_{I})$ in the diagram (4.1) induces a canonical map $\xi_{I,I^{\prime}}:\boldsymbol{\sigma}_{I^{\prime}}\hookrightarrow\overline{\boldsymbol{\sigma}_{I}}$ which we will call the adjacent map.

### v4

@@ 2.3.3. Examples for projective completions
Though the construction of the CMCX $\mathcal{T}$ and the completion $\hat{B}$ of $B$ given by Theorem 2.8 are in general unapproachable without using computer, there are still some accessible cases including the case when $B$ is a surface. In Section 7 we will review several such examples. We will show the main Theorem 2.8 gives completed period mappings in the category of projective varieties. These are cases in which the Conjecture 2.12 is true.
@@ 2.3.4. Algebraicity of weakly special subvarieties
In Hodge theory, weakly special subvarieties for a $\mathbb{Z}$-PVHS $\mathbb{V}\rightarrow B$ are analytic subvarieties of $B$ carrying special Hodge-theoretic or monodromic data. In the lanaguage of o-minimal geometry, since the period map is definable by [BKT20], weakly special subvarieties are definable analytic subvarieties of $B$ and hence must be algebraic by the definable GAGA theorem [BBT22]. This recovers and generalizes the classical result of [CDK95] that Hodge loci is a countable union of algebraic varieties.
As an application of our main theorem, we give an alternative proof of the algebraicity result on weakly special subvarieties using our main theorem 2.8. Our proof is based on the Hodge-theoretic interpretation of the completed period map $\overline{\Phi}$ provided by Theorem 2.8. We will show that weakly special subvarieties of $B$ can be extended to analytic subvarieties of $\overline{B}$ over which an extended period map $\overline{\Phi}$ exists. The boundary Hodge-theoretic data are described by the period torsors defined by Kato-Usui in [KU08, Chap. 3]. Indeed, this can be understood as a logarithmic geometry formalism of [CDK95]’s idea.
2.4. Important remark on the connectivity of strata
Similar to some related works like [Gre+23] and [GGR21], we implicitly use the assumption that every strata $Z_{I}^{*}\subset\hat{B}-B$ is connected. This is in general not the case, but we claim that we may assume it without loss of generality. Indeed, every related argument in this article only uses the partially ordered set of connected boundary strata of $\hat{B}$. In other words, we may re-index each $Z_{I}^{*}$ by its connected components $Z_{I,1}^{*},Z_{I,2}^{*},$... without essentially changing any arguments, but this will make notations in this paper unnecessarily complicated. Therefore, we decide to ignore the component index in this paper.
3. Hodge theory background
@@ 3.1. Basic settings
(3.1) $\D\Phi:B\rightarrow\Gamma\backslash$ D
where $D$ is the classifying space of polarized Hodge structures defined on the integral polarized lattice $(V_{\mathbb{Z}},Q)$ and $\check{D}$ be its compact dual. Moreover, denote $\mathbf{G}=\mathrm{Aut}(H,Q)$ and $\mathfrak{g}:=\mathrm{Lie}(G)$, then $\Gamma\leq G_{\mathbb{Z}}$ is the monodromy group of the local system $\mathbb{V}\rightarrow B$:
@@
In addition, we call $\D\Gamma\backslash$ D a Hodge variety associated to the Hodge datum $(\mathbf{G},D)$.
@@ 3.2. Nilpotent cones and nilpotent orbits
Suppose a convex polyhedral cone $\sigma\subset\mathfrak{g}_{\mathbb{Q}}$ is the convex hull of commuting nilpotent elements $\{N_{1},...,N_{n}\}$. In other words,
(3.2) $\sigma:=\{\sum_{j=1}^{n}x_{j}N_{j}\ |\ x_{1},...,x_{n}\in\mathbb{Q}_{\geq 0}\}.$
@@
    $\mathrm{exp}(\sum_{j=1}^{n}z_{j}N_{j})F\in D$ for $z_{j}\in\mathbb{C}$ and $\mathfrak{Im}(z_{j})\gg 0$ for any $j$.
[…]
It is a classical result in Hodge theory that the notions of nilpotent orbits and limiting mixed Hodge structures are closely connected:
@@
[CK82] $(\sigma,F)$ gives a nilpotent orbit if and only if $(W(\sigma),F)$ is a mixed Hodge structure polarized by any $N\in\sigma^{\circ}$, where $W(\sigma)$ is the Jacobson-Morozov weight filtration associated to any $N\in\sigma^{\circ}$, appropriately shifted.
We thus see that $\tilde{A}(\sigma)$ is naturally a mixed period domain, and the space of nilpotent orbits $A(\sigma)$ is a natural quotient of it. Indeed, $A(\sigma)$ is a mixed hodge domain as in [Kli16, Thm 6.4].
@@
Any element $g\in\mathbf{G}$ acts on the set of nilpotent orbits by:
$g(\sigma,F)\rightarrow(\mathrm{Ad}_{g}\sigma,gF)\in A(\mathrm{Ad}_{g}\sigma).$
@@ 3.2.1. Schmid’s nilpotent orbit theorem
For any $b\in\hat{B}$, choose a local neighborhood around $b\in U\subset\hat{B}$ satisfying $U\cap B\cong(\Delta^{*})^{k}\times\Delta^{n-k}$ on which we may consider the local monodromy operators $\{T_{i}\}_{1\leq i\leq k}$ around each boundary divisor and their logarithms $\{N_{i}:=\log(T_{i})\}_{1\leq i\leq k}$. By picking a local lift to $D$, we may consider the $N_{i}$ as elements of $\mathfrak{g}$. One of the main ingredients is the following interpretation of Schmid’s nilpotent orbit theorem [Sch73]:
@@
Pick a local lift to $D$ as above. Then The cone $\mathrm{Conv}(\{N_{i}\}_{1\leq i\leq k})\subset\mathfrak{g}_{\mathbb{Q}}$ is a nilpotent cone, and there is a nilpotent orbit $(\sigma_{b},F_{b})$ canonically associated to $b$ and our local lift for some $F_{b}\in\check{D}$. Changing the local lift changes the nilpotent orbit by the action of the corresponding element of $\Gamma$.
@@ 3.3. Kato–Nakayama–Usui’s construction
Let $\Sigma$ be a collection of rational nilpotent cones in $\mathfrak{g}_{\mathbb{Q}}$, and $|\Sigma|$ be its support in $\mathfrak{g}_{\mathbb{Q}}$. For $\sigma\in\Sigma$, set $\Gamma(\sigma):=\Gamma\cap\mathrm{exp}({\mathbb{Q}_{\geq 0}}\sigma)$. This is a semigroup. Denote $\Gamma_{\sigma}:=\Gamma(\sigma)^{\mathrm{gp}}$ as its associated group.
@@
    $\overline{\sigma_{\mathbb{Q}}}=\mathbb{Q}_{\geq 0}\langle\mathrm{log}(\Gamma(\sigma))\rangle$ for any $\sigma\in\Sigma$.
@@
    For any $\sigma,\tau\in\Sigma$, either $\sigma^{\circ}\cap\tau^{\circ}=\emptyset$ or $\sigma=\tau$;
@@
    For any $\sigma,\tau\in\Sigma$, suppose $\sigma^{\circ}\cap\tau^{\circ}$ is not empty, and there exists $F\in\check{D}$ such that both $(\sigma,F)$ and $(\tau,F)$ are nilpotent orbits, then $\sigma=\tau$.
@@
(3.7) $D_{\Sigma}:$ $=\{(\sigma,Z)\ |\ \sigma\in\Sigma,Z\ \text{is a}\ \sigma-\text{nilpotent orbit}\}$
@@
The first main result of Kato–Nakayama–Usui’s theory is about the (partial) completion of the Hodge variety $\D\Gamma\backslash$ D:
@@
[…]
We refer readers to [KU08, Chap. 3] for the literature of logarithmic manifolds. For our purposes, it suffices to note that these are locally ringed spaces equipped with an fs log-structure. Some important properties will be reviewed in Section 8. When $D$ is Hermitian, one may choose the (weak) fan $\Sigma$ to be the rational polyhedral decompositions constructed by classical theories (e.g., [Ash+10], [Nam76]), then the Kato–Nakayama–Usui’s construction coincide with the toroidal compactifications $\overline{\Gamma\backslash D}^{\Sigma}$. In general, the space $\Gamma\backslash D_{\Sigma}$ does not admit a structure of complex varieties.
@@
[…]
@@
If the $\mathbb{Z}$-PVHS $\mathbb{V}\rightarrow B$ can be extended to a $\mathbb{Z}$-LPVHS on $B^{\prime}$ compatible with $(\Gamma,\Sigma)$, then there exists a unique morphism of logarithmic manifolds
@@
extending $\D\Phi:B\rightarrow\Gamma\backslash$ D to $B^{\prime}$. For any $b\in B^{\prime}$, $\overline{\Phi}(b)\in\Gamma\backslash D_{\Sigma}$ is given by the nilpotent orbit $(\tau_{b},F_{b})$ for some $\tau_{b}\in\Sigma$ determined by Definition 3.9.
@@
Let $U\subset B^{\prime}$ be an analytic (resp. algebraic) subspace. If $\overline{\Phi}|_{U}$ is proper, then $U\xrightarrow{\overline{\Phi}}\overline{\Phi}(U)$ is a morphism in the category of complex analytic (resp. algebraic) spaces.
@@
This theorem is in [Usu06, Sec. 5]. Though the original theorem takes the assumption that $U$ is compact (for example, take $U=B^{\prime}=\hat{B}$), the proof only uses the properness of $\overline{\Phi}|_{U}$ and hence applies to our case verbatim. ∎
@@ 4. The combinatorial monodromy complex (CMCX)
4.1. The combinatorial monodromy cone (CMC) and adjacency map
For any $I\in\mathcal{I}$, we may choose any point $b\in Z_{I}^{*}$, a neighborhood $b\in E_{b}\subset\hat{B}$ such that $E_{b}^{*}:=E_{b}\cap B$ is isomorphic to $(\Delta^{*})^{|I|}\times\Delta^{\mathrm{dim}(B)-|I|}$, and a nearby point $b^{\prime}\in E_{b}^{*}$. For every $i\in I$ there is a monodromy logarithm $N_{i}\in\mathrm{End}(V_{b^{\prime},\mathbb{Q}})\simeq\mathfrak{g}_{\mathbb{Q}}$ around the divisor $Z_{i}$. Let $\sigma_{I,b^{\prime}}:=\mathrm{Conv}\{N_{i},i\in I\}$. There exists a linear map:
@@ 4.1. The combinatorial monodromy cone (CMC) and adjacency map
It is clear that the monodromy cone $\sigma_{I,b^{\prime}}$ is isomorphic to the image of $\mathbb{Q}_{\geq 0}^{|I|}$ under the quotient map $\mathbb{Q}^{|I|}\rightarrow\mathbb{Q}^{|I|}/\mathrm{Ker}(\lambda_{I})$. Its combinatorial structure is determined by the index set $I\subset\mathcal{I}$ and $\mathrm{Ker}(\lambda_{I})$. It does not depend on the choice of $b\in Z_{I}^{*}$, the neighborhood $E_{b}$ around $b$, or the nearby point $b^{\prime}$.
@@
The combinatorial monodromy cone (CMC) $\boldsymbol{\sigma}_{I}$ is defined to be the image of $\mathbb{Q}_{\geq 0}^{|I|}$ under the map $\mathbb{Q}^{|I|}\rightarrow\mathbb{Q}^{|I|}/\mathrm{Ker}(\lambda_{I})$.
@@
Starting from this point, for every $I\in\mathcal{I}$ we will denote $\sigma_{I}$ or $\tau_{I}$ as the original monodromy cones as a subset of some $\mathrm{End}(V_{b,\mathbb{Q}})$, while the corresponding boldface characters $\boldsymbol{\sigma}_{I}$ or $\boldsymbol{\tau}_{I}$ as the corresponding CMC’s. Note that as combinatorial cones, $\sigma_{I}\simeq\boldsymbol{\sigma}_{I}$.
For any sub-index $I^{\prime}\subset I$, the CMC $\boldsymbol{\sigma}_{I^{\prime}}$ is determined by the space $\mathrm{Ker}(\lambda_{I^{\prime}})$, and $\boldsymbol{\sigma}_{I^{\prime}}\simeq\mathrm{Conv}\{N_{i},i\in{I^{\prime}}\}\subset\mathrm{End}(V_{b^{\prime},\mathbb{Q}})$. The set-theoretic inclusion $\{\epsilon_{i},i\in{I^{\prime}}\}\hookrightarrow\{\epsilon_{i},i\in I\}$ induces $\mathbb{Q}^{|{I^{\prime}}|}\hookrightarrow\mathbb{Q}^{|I|}$, and $\mathrm{Ker}(\lambda_{I^{\prime}})=\mathrm{Ker}(\lambda_{I})\cap\mathbb{Q}^{|{I^{\prime}}|}$. We have the commutative diagram:
@@
The map $\mathbb{Q}^{|{I^{\prime}}|}/\mathrm{Ker}(\lambda_{J})\hookrightarrow\mathbb{Q}^{|I|}/\mathrm{Ker}(\lambda_{I})$ in the diagram (4.1) induces a canonical map $\xi_{I,I^{\prime}}:\boldsymbol{\sigma}_{I^{\prime}}\hookrightarrow\boldsymbol{\sigma}_{I}$ which we will call the adjacency map.